# Nonclassical dynamics of Néel vector and magnetization accompanied by THz and high-harmonic radiation from ultrafast-light-driven NiO antiferromagnet insulator

### v2

@@ Abstract
The same NiO material driven by subgap fsLP, but in combination with heavy metal (HM) layer like Pt that introduces effects [8] due to strong spin-orbit coupling (SOC), has been recently explored [9] as potential spintronic THz emitter [10, 11, 12, 13, 14, 15]. Isolated ferromagnetic metal (FM) layers (such as Co,Fe,Ni) or FM/HM bilayers have been intensely studied for nearly 30 years in order to understand ultrafast demagnetization [16, 17], as well as THz emission by these systems, which is much stronger in the case of bilayers [10, 11, 12, 13, 14, 18] than in the case of a single FM layer [19, 13, 18]. The fsLP in spintronic experiments typically has a central wavelength of nm, so that its photons have energy centered around eV which is subgap [9, 20] with respect to the separation eV between the two Hubbard bands of NiO. The spintronic phenomena in such experiments on subgap-fsLP-excited NiO have been interpreted [9] by borrowing the standard intuitive picture [10, 11, 12, 13, 14, 18] (for its recent modifications, however, via microscopic theory see Refs. [21, 22, 23]) developed for FM/HM bilayers. That is, an ultrafast spin current is somehow generated that flows from NiO into the HM layer, so that latter can convert it into charge current via the inverse spin Hall effect [24]. The time-dependent charge current is needed to obtain sizable electromagnetic (EM) radiation in the far-field (FF) region [Fig. 1], as well as to interpret [9] the enhancement [13] of the emitted radiation when switching from FM to FM/HM systems. This is because magnetic dipole radiation [19, 13, 18] from time-dependent magnetization is smaller [21] than radiation by a time-dependent charge current [Eqs. (5) vs. (6)]. However, this picture does not explain the microscopic mechanism that generates the assumed spin current in NiO/HM bilayers (only speculation exist thus far [25]), or why the frequency spectrum of such current has features within 0.1–3 THz range that is imprinted in the EM radiation detected experimentally [9]. It is obvious that fsLP will drive electrons into dynamics at its own center frequency , as well as at integer (typically odd [26, 27]) multiples of for sufficiently intense fsLP. Such high-harmonic generation (HHG) in current and EM radiation has been vigorously explored in recent years in diverse quantum materials [28, 29], including strongly correlated ones [6, 7, 30, 31, 26]. Finally, it remains unclear what kind of dynamics is obeyed by the Néel vector (as the difference of two sublattices magnetizations) and magnetization [as the sum of sublattice magnetizations, which is necessarily a nonequilibrium quantity because ], when compared to standard demagnetization [16, 17, 13] of FM layers, where magnetization vector shrinks. In the case of thin FM layers, a rapid and straightforward analysis of the direction of and its magnitude is achieved via [13] magneto-optical Kerr or Faraday effects. In contrast, they do not apply to AFs, so novel ideas have been explored [32] to detect the presumed [33] rotation of the Néel vector.
@@
In this Letter, we aim to capture all the essential physics of strongly correlated electrons within NiO, in the presence of fsLP and SOC, by employing a two-orbital Hubbard-Hund-Heisenberg (2HHH) model [2] and by simulating its time evolution via numerically exact quantum-many body methods. Its ground AF state is formed by local spins [55] at each site, that are composed of two elemental spins located on two orbitals at that site [Fig. 1], with Hund interaction included between them in order to achieve fixed and stable per atom [56, 57]. In addition, to take into account adjacent nonmagnetic HM layer in THz spintronics experiments [9], we also include SOC of the Rashba type [8, 34] into the 2HHH model of Ref. [2]. Thus, the introduced SOC models proximity effects [8, 36, 37, 38] around the NiO/HM interface, that modifies the electronic structure on the NiO side due to the HM layer (NiO, in turn, modifies bands of HM layer, but we do not include HM layer explicitly due to high computational expense). This model is placed onto a ladder geometry [Fig. 1], allowing for its spectral function [52, 53, 54] to be computed via numerically (quasi)exact simulations using time-dependent density matrix renormalization group (tDMRG) [58, 59, 60, 61] applicable to quasi-one-dimensional lattices. We complement the study with an additional set of numerical simulations using massively parallel exact diagonalization (ED) methods [62] for Hubbard-type systems, implemented within the H package [63, 64]. This technique allows us to access longer times (required for THz radiation calculations) than those possible via tensor network algorithms (like tDMRG) encountering “entanglement barrier” [65, 66].
@@
Model and Methods.—The monolayer of NiO is modeled on a tight-binding ladder [Fig. 1] with two orbitals of Ni per site only, albeit O mediated interactions are also included. The same 2HHH model, but without any SOC, was used in Ref. [2] where above-gap pumping of NiO by fsLP was studied experimentally and theoretically. Our Hamiltonian, , is built on top of the model from Ref. [2] by including possible Rashba SOC [8, 34] to describe the NiO/HM bilayer used in THz spintronic experiments [9]. The local terms of account for the Hubbard and Hund physics
@@
where denotes row vector of two creation operators and is a direction-dependent matrix hopping [35] with values for horizontal (vertical) bonds. The term is of the Rashba type [34], assumed to originate from proximity to HM layer employed experimentally [9], as often found in DFT calculations [8] on magnet/HM bilayers. Realistic parameters values for NiO are taken from prior first-principles calculations for strongly correlated electrons [72], such as from DFT+U [51] and/or DFT+dynamical mean field theory [73, 74] studies: ; eV; ; and we set and for symmetry reasons [68]. Half-filling is selected by setting the chemical potential to [2]. The magnetic field at site , , as generated by, e.g., a static impurity, induces [71] Néel “checkerboard” order ( and are two NN sites) in the ground state (GS). Nevertheless, the GS at of our simulation retains nonzero entanglement entropy, as witnessed in many recent experiments even at finite temperature [75, 76], so it is not identical to unentangled Néel ket .
The fsLP is introduced via its vector potential [gray line in Figs. 3(e) and 4(e)] of amplitude , which couples to electrons in the form of a Peierls phase [77, 78] multiplying hoppings in Eqs. (3) and (4) by a factor , so and . Here is the dimensionless parameter [79] quantifying the fsLP intensity; is the lattice constant; and fs determines the width of the Gaussian envelope which is initially centered at fs. The center frequency of the fsLP is either eV in Fig. 3, corresponding to a subgap nm wavelength commonly employed in THz spintronic experiments [9, 10, 11, 12, 13, 14, 18]; or eV in Fig. 4 for above-gap fsLP [gaps are shown in Fig. 2]. The electric field of the fsLP is linearly polarized along the -direction (i.e., diagonal to the ladder) in Fig. 1. The time evolution described by on a ladder is obtained via ED [62] by means of the H package [64, 63], where the GS was found using the Lanczos algorithm and the evolution operator is computed via a Taylor expansion considering up to terms. The time step was chosen as .
@@
Results and Discussion.—The usual first take at interpreting experiments on subgap light-driven AFIs, including NiO [85, 86], invokes [85, 87, 88, 86] a direct coupling of light magnetic field (or an effective one due to inverse Faraday effect [86]) local magnetization of AFIs. This leads to classical dynamics of the Néel vector, which rotates without changing its length, in accord with phenomenological theories [33] of Landau-Lifshitz (LL) type [89]. However, limitations of this approach are often found in experiments [86, 90], which is not surprising as light-charge coupling is much stronger [43], so electrons should be explicitly included. But the picture of classical LL dynamics [33] is appealing because it is difficult to develop intuition on how lectrons of AFI, with gapped energy spectrum [Fig. 2], can absorb subgap light to affect its magnetic ordering. Ie case of AF metals, it is easy to envisage (and calculate [81]) howthat fsLP generates photocurrent of conduction electrons, which are then spin-polarized by the magnetic background and exert spin torque [81] onto local magnetization. Its dynamics follows (for weak-intensity fsLP to avoid demagnetization [16, 17]) classical LL dynamics (valid at sufficiently high temperature and for AFIs with [66]). For AFIs, there is no such shortcut and one has to handle complexities of photoexcited Hubbard model [4, 5, 6, 7, 3], which cannot [91], in general, be reduced to studying just their final outcome on magnetic order.
@@
Indeed, our quantum many-body calculations for subgap fsLP-driven NiO reveal highly nonclassical [71, 92] (i.e., outside any description by LL-type theories [33]) dynamics of the Néel vector and magnetization in Fig. 3. That is, both vectors are changing length along the -axis (orthogonal to the ladder in Fig. 1) while not rotating at all. Nonequilibrium magnetization remains zero in plain NiO [Fig. 3(b)]. However, in SO-proximitized NiO [Fig. 3(f)], which is akin to experimentally observed [93] weak ferromagnetism in photodoped Mott insulator with native SOC. The nonzero emits magnetic dipole [Eq. (6)] THz radiation [orange line in Fig. 3(h)], which can (surprisingly, when compared to FM layers [21]) surpass the contribution from [red curve in Fig. 3(h)]. Importantly, THz emission from either of the two sources is significant only only when proximity SOC is present in NiO/HM bilayer [Fig. 3(h)], in full accord with experiments [9]. Thus, our theory explains these experiments without invoking qualitative speculations [9, 25], while showing that concepts borrowed from FM/HM systems (like interlayer spin current and spin-to-charge conversion within HM [10, 11, 13, 14, 18, 22]) are not necessary for THz emission from AFI. The magnetic dipole radiation exhibits even integer HHG [Fig. 3(g)], while odd ones are expected [26] for radiation from , as dictated by symmetry-imposed selection rules of Floquet group theory [94]. Although SOC breaks inversion symmetry, odd integer HHG is preserved [95, 27]. This offers a scheme—detect even HHG in EM radiation from AFI—which directly corroborates how magnetization dynamics in magnets driven by fsLP can be much faster [96, 45] than observed in low-energy transport experiments [25].

### v3

@@ Abstract
The same NiO material driven by subgap fsLP, but in combination with heavy metal (HM) layer (like Pt,W,Ta) introducing effects [8] due to strong spin-orbit coupling (SOC), have been very recently explored [9, 10] as spintronic THz emitter [11, 12, 13, 14, 15, 16]. Isolated ferromagnetic metal (FM) layers (such as Co,Fe,Ni) or FM/HM bilayers have been intensely studied for nearly 30 years in order to understand ultrafast demagnetization [17, 18], as well as THz emission by these systems, which is much stronger in the case of bilayers [11, 12, 13, 14, 15, 19] than in the case of a single FM layer [20, 14, 19]. The fsLP in spintronic experiments typically has a central wavelength of nm, so that its photons have energy centered around eV which is subgap [9, 10, 21] with respect to the separation eV between the two Hubbard bands of NiO. The spintronic phenomena in such experiments on subgap-fsLP-excited NiO have been interpreted [9, 10] by borrowing the standard intuitive picture [11, 12, 13, 14, 15, 19] (for its recent modifications, however, via microscopic theory see Refs. [22, 23, 24]) developed for FM/HM bilayers. That is, an ultrafast spin current is somehow generated that flows from NiO into the HM layer, so that latter can convert it into charge current via the inverse spin Hall effect [25]. The time-dependent charge current is needed to obtain sizable electromagnetic (EM) radiation in the far-field (FF) region [Fig. 1], as well as to interpret [9, 10] the enhancement [14] of the emitted radiation when switching from FM to FM/HM systems. This is because magnetic dipole radiation [20, 14, 19] from time-dependent magnetization is smaller [22] than radiation by a time-dependent charge current. However, this picture does not explain the microscopic mechanism that generates the assumed spin current in NiO/HM bilayers (only speculation exist thus far [26]), or why the frequency spectrum of such current has features within 0.1–3 THz range that is imprinted in the EM radiation detected experimentally [9, 10]. It is obvious that fsLP will drive electrons into dynamics at its own center frequency , as well as at integer (typically odd [27, 28]) multiples of for sufficiently intense fsLP. Such high-harmonic generation (HHG) in current and EM radiation has been vigorously explored in recent years in diverse quantum materials [29, 30], including strongly correlated ones [6, 7, 31, 32, 27]. Finally, it remains unclear what type of dynamics is obeyed by the Néel vector (as the difference of two sublattices magnetizations) and magnetization [as the sum of sublattice magnetizations, which is necessarily a nonequilibrium quantity because ], when compared to standard demagnetization [17, 18, 14] of FM layers, where magnetization vector shrinks. In the case of thin FM layers, a rapid and straightforward analysis of the direction of and its magnitude is achieved via [14] magneto-optical Kerr or Faraday effects. In contrast, they do not apply to AFs, so novel ideas have been explored [33] to detect the presumed [34, 35, 10] rotation of the Néel vector.
@@
In this Letter, we aim to capture all the essential physics of strongly correlated electrons within NiO, in the presence of fsLP and SOC, by employing a two-orbital Hubbard-Hund-Heisenberg (2HHH) model [2] and by simulating its time evolution via numerically exact quantum-many body methods. Its ground AF state is formed by local spins [58] at each site, that are composed of two elemental spins located on two orbitals at that site [Fig. 1], with Hund interaction included between them in order to achieve fixed and stable per atom [59, 60]. In addition, to take into account adjacent nonmagnetic HM layer in THz spintronics experiments [9, 10], we also include SOC of the Rashba type [8, 36] into the 2HHH model of Ref. [2]. Thus, the introduced SOC models proximity effects [8, 38, 39, 40, 41] around the NiO/HM interface, that modifies the electronic structure on the NiO side due to the HM layer (NiO, in turn, modifies bands of HM layer, but we do not include HM layer explicitly due to high computational expense). This model is placed onto a ladder geometry [Fig. 1], allowing for its spectral function [55, 56, 57] to be computed via numerically (quasi)exact simulations using time-dependent density matrix renormalization group (tDMRG) [61, 62, 63, 64] applicable to quasi-one-dimensional lattices. We complement the study with an additional set of numerical simulations using massively parallel exact diagonalization (ED) methods [65] for Hubbard-type systems, implemented within the H package [66, 67]. This technique allows us to access longer times (required for THz radiation calculations) than those possible via tensor network algorithms (like tDMRG) encountering “entanglement barrier” [68, 69].
@@
Model and Methods.—The monolayer 111We use monolayer of NiO and do not include HM layer explicitly due to high computational expense of our numerically exact calculations. Nevertheless, this limitation is not a drawback for capturing essential features of experiments as they find [10] the highest intensity of emitted THz radiation for the thinnest NiO layer driven by subgap fsLP. of NiO is modeled on a tight-binding ladder [Fig. 1] with two orbitals per site hosting Ni only, while O atom is not modeled explicitly albeit O-mediated interactions areincluded. The same 2HHH model, but without any SOC, was used in Ref. [2] where above-gap pumping of NiO by fsLP was studied experimentally and theoretically. Our Hamiltonian, , is built on top of the model from Ref. [2] by including possible Rashba SOC [8, 36] to describe the NiO/HM bilayer used in THz spintronic experiments [9, 10]. The local terms of account for the Hubbard and Hund physics
@@
where denotes row vector of two creation operators and is a direction-dependent matrix hopping [37] with values for horizontal (vertical) bonds. The term is of the Rashba type [36], assumed to originate from proximity to HM layer employed experimentally [9, 10], as found in DFT calculations on FM/HM [8] of AFI/HM [40] bilayers. Realistic parameters values for NiO are taken from prior first-principles calculations for strongly correlated electrons [76], such as from DFT+U [54] and/or DFT+dynamical mean field theory [77, 78] studies: ; eV; ; and we set and for symmetry reasons [71]. Half-filling is selected by setting the chemical potential to [2]. The magnetic field at site , , as generated by, e.g., a static impurity, induces [75] Néel “checkerboard” order ( and are two NN sites) in the ground state (GS). Nevertheless, the GS in our simulations retains nonzero entanglement, as witnessed in many recent experiments even at finite temperature [79, 80], so it is not identical to unentangled Néel ket .
The fsLP is introduced via its vector potential [gray line in Figs. 3(e) and 4(e)] of amplitude , which couples to electrons in the form of a Peierls phase [81, 82] multiplying hoppings in Eqs. (3) and (4) by a factor , so and . Here is the dimensionless parameter [83] quantifying the fsLP intensity; is the lattice constant; and fs determines the width of the Gaussian envelope which is initially centered at fs. The center frequency of the fsLP is either eV in Fig. 3, corresponding to a subgap nm wavelength commonly employed in THz spintronic experiments [9, 11, 12, 13, 14, 15, 19, 10]; or eV in Fig. 4 for above-gap fsLP [gaps are shown in Fig. 2]. The electric field of the fsLP is linearly polarized along the -direction (i.e., diagonal to the ladder) in Fig. 1. The time evolution described by on a ladder is obtained via ED [65] by means of the H package [67, 66], where the GS was found using the Lanczos algorithm and the evolution operator is computed via a Taylor expansion considering up to terms. The time step was chosen as .
@@
Results and Discussion.—The usual first take at interpreting experiments on subgap light-driven AFIs, including NiO [89, 90, 10], invokes [89, 91, 92, 90] a direct coupling of light magnetic field (or an effective one due to inverse Faraday [90] or inverse Cotton-Mouton [10] effects) to local magnetization of AFIs. This leads to classical dynamics of the Néel vector, which rotates without changing its length [10], in accord with phenomenological theories [34, 35] of Landau-Lifshitz (LL) type [93]. However, limitations of this approach are often found in experiments [90, 94], which is not surprising as light-charge coupling is much stronger [46], so electrons should be explicitly included. But the picture of classical LL dynamics [34, 35, 10] is appealing because it is difficult to develop intuition on how lectrons of AFI, with gapped energy spectrum [Fig. 2], absorb subgap light and then affect its magnetic ordering. Ie case of AF metals, it is easy to envisage (and calculate [85]) howthat fsLP generates photocurrent of conduction electrons, which are then spin-polarized by the magnetic background and exert spin torque [85] onto local magnetization. Its dynamics follows (for weak-intensity fsLP to avoid demagnetization [17, 18]) classical LL dynamics (valid at sufficiently high temperature and for AFIs with [69]). For AFIs, there is no such shortcut and one has to handle complexities of photoexcited Hubbard model [4, 5, 6, 7, 3], which cannot [95], in general, be reduced to studying just their final outcome on magnetic order.
@@
Indeed, our quantum many-body calculations for subgap fsLP-driven NiO reveal highly nonclassical [75, 96] (i.e., outside any description by LL-type theories [34, 10]) dynamics of the Néel vector and magnetization in Fig. 3. That is, both vectors are changing length along the -axis (orthogonal to the ladder in Fig. 1) while not rotating at all. Nonequilibrium magnetization remains zero in plain NiO [Fig. 3(b)]. However, in SO-proximitized NiO [Fig. 3(f)], which is akin to experimentally observed [97] weak ferromagnetism in photodoped Mott insulator with native SOC. The nonzero emits magnetic dipole [Eq. (6)] THz radiation [orange line in Fig. 3(h)], which can (surprisingly, when compared to FM layers [22]) surpass the contribution from [red curve in Fig. 3(h)]. Importantly, THz emission from both of these two sources is significant only only when proximity SOC is present in NiO/HM bilayer [Fig. 3(h)], in full accord with experiments [9, 10]. Thus, our theory explains these experiments without invoking qualitative speculations [9, 26, 10], while showing that concepts borrowed from FM/HM systems (like interlayer spin current and spin-to-charge conversion within HM [11, 12, 14, 15, 19, 23]) are not necessary for THz emission from AFI. The magnetic dipole radiation exhibits even integer HHG [Fig. 3(g)], while odd ones are expected [27] for radiation from , as dictated by symmetry-imposed selection rules of Floquet group theory [98]. Although SOC breaks inversion symmetry, odd integer HHG is preserved [99, 28]. This offers a scheme—detect even HHG in EM radiation from AFI—which directly corroborates how magnetization dynamics in magnets driven by fsLP can be much faster [100, 48] than observed in low-energy transport experiments [26].
@@ Acknowledgements.
- Rongione et al. [2023] E. Rongione, O. Gueckstock, M. Mattern, O. Gomonay, H. Meer, C. Schmitt, R. Ramos, T. Kikkawa, M. Mičica, E. Saitoh, et al., Emission of coherent thz magnons in an antiferromagnetic insulator triggered by ultrafast spin–phonon interactions, Nat. Commun. 14, 1818 (2023).
@@
- Gomonay and Loktev [2010] H. V. Gomonay and V. M. Loktev, Spin transfer and current-induced switching in antiferromagnets, Phys. Rev. B 81, 144427 (2010).
@@
- Dolui et al. [2022] K. Dolui, A. Suresh, and B. K. Nikolić, Spin pumping from antiferromagnetic insulator spin-orbit-proximitized by adjacent heavy metal: a first-principles Floquet-nonequilibrium Green function study, J. Phys. Mater. 5, 034002 (2022).
@@
- Note [1] We use monolayer of NiO and do not include HM layer explicitly due to high computational expense of our numerically exact calculations. Nevertheless, this limitation is not a drawback for capturing essential features of experiments as they find [10] the highest intensity of emitted THz radiation for the thinnest NiO layer driven by subgap fsLP.